OpenStudy (anonymous):

If f(x) = 2x and g(x) = x-4, what is the value of f[g(3)]?

OpenStudy (anonymous):

f composed with g = 2(x-4) then evaluate that @ x=3, so -2.

OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = (x - 3)/2, what is the value of f[g(-5)]? would this be -7 then

OpenStudy (anonymous):

f(g(x))= 2(2x-3)/2 + 3 , the two's cancelled out it becomes 2x-3+3= 2x replace x by -5 you'll get -10